# Learning Group Invariant Calabi–Yau Metrics by Fundamental Domain Projections

### v1

@@ Abstract
We present new invariant machine learning models that approximate the Ricci-flat metric on Calabi–Yau (CY) manifolds with discrete symmetries. We accomplish this by combining the $\phi$-model of the cymetric package with non-trainable, $G$-invariant, canonicalization layers that project the $\phi$-model’s input data (i.e. points sampled from the CY geometry) to the fundamental domain of a given symmetry group $G$. These $G$-invariant layers are easy to concatenate, provided one compatibility condition is fulfilled, and combine well with spectral $\phi$-models. Through experiments on different CY geometries, we find that, for fixed point sample size and training time, canonicalized models give slightly more accurate metric approximations than the standard $\phi$-model. The method may also be used to compute Ricci-flat metric on smooth CY quotients. We demonstrate this aspect by experiments on a smooth $\mathbb{Z}^{2}_{5}$ quotient of a 5-parameter quintic CY manifold.
@@ 1 Introduction
Admittedly, the accuracy of ML algorithms has been observed to improve on more symmetric spaces [18, 19, 24, 27], and the training time (for similar performance) is slightly longer on non-symmetric spaces [27]. These differences, which are small compared to the effects seen in pre-ML methods, indicate that the ML models learn the symmetries of the data they are trained on. Symmetries have also been used to test the accuracy of ML predictions of CY metrics [17]. However, discrete symmetries have, so far, not played any significant role in the design of ML models for CY metrics. In this paper, we will perform a first exploration of this topic, using novel methods to construct group invariant ML models. [^7] In methods based on spectral basis, such as [16, 23] and [18, 19], one can restrict to group invariant elements of the basis; to the knowledge of the authors this has not been explored to any greater degree.
The motivation for our study is two-fold. First, as we will review below, in the ML literature there is ample evidence that for symmetric data, equivariant and invariant models outperform networks that are agnostic about symmetries in the data. The Geometric Deep Learning program [28, 29] shows how symmetries explain the success of state-of-the-art ML models such as such as CNNs, GNNs, and Transformers; ref. [30] analyse the CNNs using the concept of covariance in theoretical physics. Formally, it has been proven that group invariant neural networks are universal approximators for group invariant functions [31], thus symmetry-enforcing ML models are as expressive as the symmetry-agnostic ones.
Secondly, from the perspective of string theory, there is phenomenological motivation to study non-simply connected CY geometries, obtained by quotienting a CY manifold by a freely acting discrete symmetry $G$. $G$-invariant ML models, provides a new way to compute the Ricci-flat metrics of such quotients. This problem has been explored using Donaldson’s algorithm in the past [8], with positive effects on the accuracy of the metric approximation. Ref. [8] also explains the technical subtleties associated with the creation of a group invariant spectral basis; this provides further motivation to explore invariant ML methods, as do recent application of ML methods to the computations of heterotic Yukawa couplings on CY manifolds with discrete symmetries [32, 33].
@@
For our studies, we will explore CY geometries created as complete intersections in complex projective spaces, so-called CICYs [34]. We will use the cymetric package, introduced in [21, 22], which allows us to sample points on such spaces, and provide several ML model architectures that either predict the function $\phi$, or the metric directly. In this work, we will limit our attention to the $\phi$-model. To create group invariant $\phi$-models, we will include linear, non-trainable layers into the neural network architecture. These layers pre-process the the input data of the $\phi$-model, and ensure that it is $G$-invariant by projecting to a fundamental domain of the group $G$. This idea, which is often referred to as canonicalization, has recently been proposed in [35, 36]. Our main contribution is a comparison of the performance and accuracy of the canonicalized ML models with the standard $\phi$-model. We also discuss the similarities and differences of the canonicalization approach with the spectral networks suggested by [24], and how canonicalization can be combined with spectral $\phi$-models. Finally, we study how canonicalized ML models can be used to compute the metric on a non-simply connected CY manifold.
This paper is organised as follows. In section 2, we briefly review related work on machine learned, Ricci-flat CY metrics, and outline the key approaches to group invariant ML models. Section 3 describes CY geometry and the ML model we will use for our experiments. We also recall how the spectral layer is constructed. In section 4, we present our main contribution: how to construct canonicalization layers for the symmetry groups relevant for CY point samples; experiments using these techniques are presented and discussed in section 5. Some technical aspects of our analysis are described in appendices. In section 6 we apply canonicalization layers to the computation of Ricci-flat CY metrics on smooth quotients of a 5-parameter quintic CY. Conclusions and discussions are given in section 7.
@@ 2 Related work
To date, three open-source ML libraries have been developed. While we will only explore one of these libraries in this work, the key features of the other packages may also be of interest to the reader, and are therefor given here
@@
There are different approaches to the construction of invariant machine learning models. Following Yarotsky [31], see also [35], one may distinguish between symmetrization-based and intrinsic approaches. Following the first approach, one may, for example, construct $G$-invariant models by augmenting the training data by $G$-transformed copies with the same label as the original data point. The network will then learn that data points related by $G$-transformations give the same output, as required by invariance. When applied to unseen data, a network trained on augmented data will perform better than a network trained on non-augmented data. As discussed in ref. [41], from a mathematical point of view, and under exact invariance, data augmentation is equivalent to averaging over orbits of the group $G$; it can thus be proven that this reduces the variance of the prediction of a model.
Intrinsic approaches instead produce invariant networks by imposing restrictions on the network design. CNNs, which are equivariant under translations, are early incarnations of this idea [42, 43]. Other network designs that are invariant under other group transformations include [44, 45, 46, 47, 48]; these and related ideas are further discussed in [29] and the review [49]. Common to these designs is that in- or equivariance is implemented by weight-sharing in some layer of the network.
Restricting the trainable weights of some layer in the network is not the only intrinsic way to achieve invariant/equivariant layers. Yarotsky has shown, using classical polynomial invariant theories, that an extra polynomial layer wich is invariant/equivariant and complete, results in an invariant/equivariant network [31]. This is particularly useful for finite groups, where complete invariant layers can be constructed from a finite number of polynomial invariants. For continuous groups, invariant layers can also be complete, but, lacking completeness, one cannot prove that they are universal invariant approximators [31]. The spectral layer [24] (described briefly above and in detail below) is, up to rescaling by the norm, an incarnation of such an invariant layer for the continuous group $\mathbb{C}^{*}$.
@@ 3 Machine learning Ricci-flat CY metrics
In this section, we review some aspects of the construction of examples of CY manifolds, and present some facts about their symmetries. A more detailed discussion of the symmetries is relegated to appendix A. We then describe the key points of the cymetric package, and the ML methods we employ to learn CY metrics. For more details on the cymetric package, we refer to [21, 22].
@@ 3.1 CY construction
For our experiments, we need explicit realizations of CY geometries. In this paper, we will restrict our studies to the Complete Intersection Calabi Yau (CICY) manifolds [34, 50]. [^9] It would be interesting to extend our studies to CY manifolds in toric ambient spaces. The cymetric package has point generators also for such spaces, and the methods we develop here can be adapted to this setting. We leave this generalisation for future work. These are constructed as the zero loci of a set of homogeneous polynomials $p_{i}$ in an ambient space $\mathbb{CP}^{n}_{1}\times...\times\mathbb{CP}^{n}_{2}$. A prime example is the Fermat quintic, defined by the solutions of
@@
We want to learn the Ricci-flat metric on some CICY manifold $X$ with dimension $n$. Starting from some reference Kähler form $J^{\prime}$, we know, by the CY theorem, that there exists a unique Kähler form $J$, in the same cohomology class as $J^{\prime}$, which gives a Ricci-flat metric.
@@ 3.2 The $\phi$-model of the cymetric package
where $\kappa$ is some complex constant and $\Omega$ is the no-where vanishing holomorphic $(n,0)$-from. [^12] Here $J$ is predicted by the network and $\kappa$ and $\Omega$ are computed by the point generator. With this at hand, the first loss function is defined as
@@
In the cymetric package, three more constraints are encoded as custom loss functions $\mathcal{L}_{i}$: the Kähler constraint $\mathrm{d}J=0$ gives $\mathcal{L}_{\text{dJ}}$, preserving the Kähler class is enforced by $\mathcal{L}_{\text{Kclass}}$, [^14] The Kähler class loss function is necessary when learning the metric on any CY manifold with $h^{(1,1)}>1$. It reduces to a volume-preserving constraint when $h^{(1,1)}=1$. The cymetric routines require such a loss since they can be applied to any CICY manifold, and CY manifolds defined as hypersurface in the toric ambient spaces contained in the Kreuzer–Skarke list [55].
@@
Here $\alpha_{i}$ are tuneable hyperparameters of the model that are set by the user. In our experiments, we use the default values for the cymetric losses in the $\phi$-model; thus the Kähler and Ricci losses are disabled and $\alpha_{i}=1$. We refer the reader to [21, 22] for more discussions of the loss functions.
@@ 3.3 Spectral layers
It may nevertheless appear surprising that such feature engineering is needed. The homogeneous rescaling is not a symmetry of the CY manifold; this is a complex, $n$-dimensional space with $n$ holomorphic coordinates. However, as we have described in the previous section, the input data of the cymetric models is not an $n$-dimensional complex tuple. Rather, for a CICY with ambient space $\mathbb{CP}^{n_{1}}\times..\times\mathbb{CP}^{n_{k}}$, the point generators of cymetric produces $\sum_{i}2(n_{i}+1)$-dimensional real input data tuples (with entries in the range $[0,1]$). In this pre-processing the points on the CY space are separated into coordinate patches (that overlap but for zero measure sets). The training is then done locally, using the transition loss function to impose that the function matches on patch overlaps. It is an empirical fact that the transition loss function is around $10^{-3}$ for standard $\phi$-models (trained for 10 epochs on the Fermat quintic). As we will see in section 5, with the addition of a spectral layer in the network, the transition loss function is of the order $10^{-9}$, which indicates the prediction of $\phi$ is indeed a function.
@@
Figure 1: A dense neural network with an invariant layer. The invariant layer does not have trainable parameters. It’s sole purpose is to put the input data, $z_{i}$, in a form, $\alpha_{i}(z)$ that is invariant under the relevant symmetry group.
@@ 4 Invariant metric by fundamental domain projections
for some group $G$ acting on $X$, one can construct a $G$-invariant neural network by
@@
It is important to note that the projection into a fundamental domain can give rise to discontinuities. For a given fundamental domain $\mathcal{F}$ of some group $G$ acting on some manifold $X$, there exist points on $X$ which will get mapped to the boundary of $\mathcal{F}$. Moving infinitesimally within $X$ can then result into discontinuous jumps in $\mathcal{F}$. This can lead to problems computing global properties like the volume which require integration over the entire manifold. We will talk about this further in section 5 and the conclusion.
@@ 4.1 Examples
In the following examples we will consider symmetries on the Fermat quintic, whose defining equation was given in eq. (3). Let $f\circ h:X\to\mathbb{R}$ be a neural network learning the function $\phi(z)$ in the Ricci-flat metric $g_{\textrm{CY}}(z)=g_{\textrm{FS}}(z)+\partial\overline{\partial}\phi(z)$. Let $\mathbf{z}\in\mathbb{C}^{5}$ encode $z$ as a vector and let $\rho:G\to\textrm{GL}(5,\mathbb{C})$ be a representation of some symmetry group $G$. A $G$-invariant network is then constructed by defining $h(\mathbf{z})=\rho(\pi(z))\mathbf{z}$ together with an appropriate choice of $\pi(z)$. In the following, we will construct examples considering several different symmetry groups $G$. In each example, we specify explicitly the action of $G$ on the ambient space, the fundamental domain $\mathcal{F}$ and the projection map $h$ on $\mathcal{F}$.
@@ 4.1.1 Homogeneous rescalings
In case more than one coordinate has the biggest norm, then we choose to divide by the coordinate with smallest index. [^17] This is straightforwardly implemented in TensorFlow— by finding the largest coordinate with tf.math.argmax(tf.abs(complex_input))— and then dividing the input by said coordinate.
@@
In case of a product of complex projective spaces, we apply homogeneous rescaling as described above for the coordinates of each factor, separately.
@@ 4.1.2 Permutations
Let $G=S_{5}$, which acts by permutation of variables. Clearly, the action of this group allows to reorder the entries of a tuple $(z_{0},...z_{n})$ in any way we desire, and we may select a fundamental domain by any consistent order. Thus, given a set of points $[z_{0}:z_{1}:z_{2}:z_{3}:z_{4}]$, we define the fundamental domain
@@
where we have used Cauchy’s two line notation for $\pi(z)\in S_{5}$. The first line corresponds to indices $i$ and the second line to their permutations $\sigma(i)$. In terms of $\mathbf{z}\in\mathbb{C}^{5}$, we have

### v2

@@ Abstract
We present new invariant machine learning models that approximate the Ricci-flat metric on Calabi–Yau (CY) manifolds with discrete symmetries. We accomplish this by combining the $\phi$-model of the cymetric package with non-trainable, $G$-invariant, canonicalization layers that project the $\phi$-model’s input data (i.e. points sampled from the CY geometry) to the fundamental domain of a given symmetry group $G$. These $G$-invariant layers are easy to concatenate, provided one compatibility condition is fulfilled, and combine well with spectral $\phi$-models. Through experiments on different CY geometries, we find that, for fixed point sample size and training time, canonicalized models give slightly more accurate metric approximations than the standard $\phi$-model. The method may also be used to the compute Ricci-flat metric on smooth CY quotients. We demonstrate this aspect by experiments on a smooth $\mathbb{Z}^{2}_{5}$ quotient of a 5-parameter quintic CY manifold.
@@ 1 Introduction
Admittedly, the accuracy of ML algorithms has been observed to improve on more symmetric spaces [18, 19, 24, 27], and the training time (for similar performance) is slightly longer on non-symmetric spaces [27]. These differences, which are small compared to the effects seen in pre-ML methods, indicate that the ML models learn the symmetries of the data they are trained on. Symmetries have also been used to test the accuracy of ML predictions of CY metrics [17]. However, discrete symmetries have, so far, not played any significant role in the design of ML models for CY metrics. In this paper, we will perform a first exploration of this topic, using novel methods to construct group invariant ML models. [^7] In ML methods based on spectral basis, such as [16, 23] and [18, 19], one can restrict to group invariant elements of the basis; to the knowledge of the authors this has not been explored to any greater degree.
The motivation for our study is two-fold. First, as we will review below, in the ML literature there is ample evidence that for symmetric data, equivariant and invariant models outperform networks that are agnostic about symmetries in the data. The Geometric Deep Learning program [28, 29] shows how symmetries explain the success of state-of-the-art ML models such as CNNs, GNNs, and Transformers; ref. [30] analyse the CNNs using the concept of covariance in theoretical physics. Formally, it has been proven that group invariant neural networks are universal approximators for group invariant functions [31], thus symmetry-enforcing ML models are as expressive as the symmetry-agnostic ones.
Secondly, from the perspective of string theory, there is phenomenological motivation to study non-simply connected CY geometries, obtained by quotienting a CY manifold by a freely acting discrete symmetry $G$. $G$-invariant ML models provide a new way to compute the Ricci-flat metrics of such quotients. This problem has been explored using Donaldson’s algorithm in the past [8], with positive effects on the accuracy of the metric approximation. Ref. [8] also explains the technical subtleties associated with the creation of a group invariant spectral basis; this provides further motivation to explore invariant ML methods, as do recent application of ML methods to the computations of heterotic Yukawa couplings on CY manifolds with discrete symmetries [32, 33].
@@
For our studies, we will explore CY geometries created as complete intersections in complex projective spaces, so-called CICYs [34]. We will use the cymetric package, introduced in [21, 22], which allows us to sample points on such spaces, and provide several ML model architectures that either predict the function $\phi$, or the metric directly. In this work, we will limit our attention to the $\phi$-model. To create group invariant $\phi$-models, we will include non-trainable layers into the neural network architecture. These layers pre-process the the input data of the $\phi$-model, and ensure that it is $G$-invariant by projecting to a fundamental domain of the group $G$. This idea, which is often referred to as canonicalization, has recently been proposed in [35, 36]. Our main contribution is a comparison of the performance and accuracy of the canonicalized ML models with the standard $\phi$-model. We also discuss the similarities and differences of the canonicalization approach with the spectral networks suggested by [24], and how canonicalization can be combined with spectral $\phi$-models. Finally, we study how canonicalized ML models can be used to compute the metric on a non-simply connected CY manifold. [^8] The code for the canonicalization layers can be found in https://github.com/jake997/invariant_layers_cymetric.
This paper is organised as follows. In section 2, we briefly review related work on machine learned, Ricci-flat CY metrics, and outline the key approaches to group invariant ML models. Section 3 describes CY geometry and the ML model we will use for our experiments. We also recall how the spectral layer is constructed. In section 4, we present our main contribution: how to construct canonicalization layers for the symmetry groups relevant for CY point samples; experiments using these techniques are presented and discussed in section 5. Some technical aspects of our analysis are described in appendices. In section 6, we apply canonicalization layers to the computation of Ricci-flat CY metrics on smooth quotients of a 5-parameter quintic CY. Conclusions and discussions are given in section 7.
@@ 2 Related work
To date, three open-source ML libraries have been developed. While we will only explore one of these libraries, the key features of the other packages may also be of interest to the reader, and are therefor given here
@@
There are different approaches to the construction of invariant machine learning models. Following Yarotsky [31], see also [35], one may distinguish between symmetrization-based and intrinsic approaches. In the first approach, one may, for example, construct $G$-invariant models by augmenting the training data by $G$-transformed copies with the same label as the original data point. The network will then learn that data points related by $G$-transformations give the same output, as required by invariance. When applied to unseen data, a network trained on augmented data will perform better than a network trained on non-augmented data. As discussed in ref. [41], from a mathematical point of view, and under exact invariance, data augmentation is equivalent to averaging over orbits of the group $G$; it can thus be proven that this reduces the variance of the prediction of a model.
Intrinsic approaches instead produce invariant networks by imposing restrictions on the network design. CNNs, which are equivariant under translations, are early incarnations of this idea [42, 43]. Other network designs that are invariant under group transformations include [44, 45, 46, 47, 48]; these and related ideas are further discussed in [29] and the review [49]. Common to these designs is that in- or equivariance is implemented by weight-sharing in some layer of the network.
Restricting the trainable weights of some layer in the network is not the only intrinsic way to achieve invariant/equivariant layers. Yarotsky has shown, using classical polynomial invariant theory, that an extra polynomial layer which is invariant/equivariant and complete, results in an invariant/equivariant network [31]. This is particularly useful for finite groups, where complete invariant layers can be constructed from a finite number of polynomial invariants. For continuous groups, invariant polynomial layers can also be constructed, but, lacking completeness, one cannot prove that they are universal invariant approximators [31]. The spectral layer [24] (described briefly above and in detail below) is, up to rescaling by the norm, an incarnation of such an invariant layer for the continuous group $\mathbb{C}^{*}$.
@@ 3 Machine learning Ricci-flat CY metrics
In this section, we review some aspects of the construction of examples of CY manifolds, and their symmetries. A more detailed discussion of the symmetries is relegated to appendix A. We then describe the key points of the cymetric package, and the ML methods we employ to learn CY metrics. For more details on the cymetric package, we refer to [21, 22].
@@ 3.1 CY construction
For our experiments, we need explicit realizations of CY geometries. In this paper, we will restrict our studies to the Complete Intersection Calabi-Yau (CICY) manifolds [34, 50]. [^10] It would be interesting to extend our studies to CY manifolds in toric ambient spaces. The cymetric package has point generators also for such spaces, and the methods we develop here can be adapted to this setting. We leave this generalisation for future work. These are constructed as the zero loci of a set of homogeneous polynomials $p_{i}$, $i=0,\dots,N$, in an ambient space $\mathbb{CP}^{n_{1}}\times...\times\mathbb{CP}^{n_{N}}$, such that $\dim(CY)=n_{1}+\dots+n_{N}-N$. A prime example is the Fermat quintic, defined by the solutions of
@@
We want to learn the Ricci-flat metric on some CICY manifold $X$ with dimension $n$. Starting from a reference Kähler form $J^{\prime}$, we know, by the CY theorem, that there exists a unique Kähler form $J$, in the same cohomology class as $J^{\prime}$, which gives a Ricci-flat metric.
@@ 3.2 The $\phi$-model of the cymetric package
where $\kappa$ is some complex constant and $\Omega$ is the no-where vanishing holomorphic $(3,0)$-from. [^13] Here $J$ is predicted by the network and $\kappa$ and $\Omega$ are computed by the point generator. With this at hand, the first loss function is defined as
@@
In the cymetric package, three more constraints are encoded as custom loss functions $\mathcal{L}_{i}$: the Kähler constraint $\mathrm{d}J=0$ gives $\mathcal{L}_{\text{dJ}}$, preserving the Kähler class is enforced by $\mathcal{L}_{\text{Kclass}}$ [^15] The Kähler class loss function is necessary when learning the metric on any CY manifold with $h^{(1,1)}>1$. It reduces to a volume-preserving constraint when $h^{(1,1)}=1$. The cymetric routines require such a loss since they can be applied to any CICY manifold, and CY manifolds defined as hypersurfaces in the toric ambient spaces contained in the Kreuzer–Skarke list [55]. and the vanishing Ricci-scalar is given by $\mathcal{L}_{\text{Ricci}}$. The total loss function is thus given by
@@
Here, $\alpha_{i}$ are tuneable hyperparameters of the model that are set by the user. In our experiments, we use the default values for the cymetric losses in the $\phi$-model; thus the Kähler and Ricci losses are disabled and $\alpha_{i}=1$. We refer the reader to [21, 22] for more discussions of the loss functions.
@@ 3.3 Spectral layers
It may nevertheless appear surprising that such feature engineering is needed. The homogeneous rescaling is not a symmetry of the CY manifold; this is a complex, $n$-dimensional space with $n$ holomorphic coordinates. However, as we have described in the previous section, the input data of the cymetric models is not an $n$-dimensional complex tuple. Rather, for a CICY with ambient space $\mathbb{CP}^{n_{1}}\times..\times\mathbb{CP}^{n_{k}}$, the point generators of cymetric produces $\sum_{i}2(n_{i}+2)$-dimensional real input data tuples (with entries in the range $[0,1]$). In this pre-processing the points on the CY space are separated into coordinate patches (that overlap but for zero measure sets). The training is then done locally, using the transition loss function to impose that the function matches on patch overlaps. It is an empirical fact that the transition loss function is around $10^{-3}$ for standard $\phi$-models (trained for 10 epochs on the Fermat quintic). As we will see in section 5, with the addition of a spectral layer in the network, the transition loss function is of the order $10^{-9}$ already at initialization of the network, which indicates the prediction of $\phi$ is indeed a function.
@@
Figure 1: A dense neural network with an invariant layer. The invariant layer does not have trainable parameters. Its sole purpose is to put the input data, $z_{i}$, in a form, $\alpha_{i}(z)$ that is invariant under the relevant symmetry group.
@@ 4 Invariant metric by fundamental domain projections
for a group $G$ acting on $X$, one can construct a $G$-invariant neural network by
@@
It is important to note that the projection into a fundamental domain can give rise to discontinuities. For a given fundamental domain $\mathcal{F}$ of some group $G$ acting on some manifold $X$, there exist points on $X$ which will get mapped to the boundary of $\mathcal{F}$. Moving infinitesimally within $X$ can then result in discontinuous jumps in $\mathcal{F}$. For the metric approximations we construct, this can lead to problems computing global properties, like the volume which require integration over the entire manifold. We will talk about this further in section 5 and the conclusion.
@@ 4.1 Examples
In the following examples, we will consider symmetries on the Fermat quintic, whose defining equation was given in eq. (3). Let $f\circ h:X\to\mathbb{R}$ be a neural network learning the function $\phi(z)$ in the Ricci-flat metric $g_{\textrm{CY}}(z)=g_{\textrm{FS}}(z)+\partial\overline{\partial}\phi(z)$. Let $\mathbf{z}\in\mathbb{C}^{5}$ encode $z$ as a vector and let $\rho:G\to\textrm{GL}(5,\mathbb{C})$ be a representation of some symmetry group $G$. A $G$-invariant network is then constructed by defining $h(\mathbf{z})=\rho(\pi(z))\mathbf{z}$ together with an appropriate choice of $\pi(z)$. In the following, we will construct examples considering several different symmetry groups $G$. In each example, we specify explicitly the action of $G$ on the ambient space, the fundamental domain $\mathcal{F}$ and the projection map $h$ on $\mathcal{F}$.
@@ 4.1.1 Homogeneous rescalings
In case more than one coordinate has the biggest norm, we choose to divide by the coordinate with smallest index. [^18] This is straightforwardly implemented in TensorFlow— by finding the largest coordinate with tf.math.argmax(tf.abs(complex_input))— and then dividing the input by said coordinate.
@@
When the ambient space is a product of complex projective spaces, we apply homogeneous rescaling as described above for the coordinates of each factor, separately.
@@ 4.1.2 Permutations
Let $G=S_{5}$, which acts by permutation of variables. Clearly, the action of this group allows to reorder the entries of a tuple $(z_{0},...,z_{n})$ in any way we desire, and we may select a fundamental domain by any consistent order. Thus, given a set of points $[z_{0}:z_{1}:z_{2}:z_{3}:z_{4}]$, we define the fundamental domain
@@
where we have used Cauchy’s two line notation for $\pi(z)\in S_{5}$, where the first line corresponds to indices $i$ and the second line to their permutations $\sigma(i)$. In terms of $\mathbf{z}\in\mathbb{C}^{5}$, we have